OpenStudy (anonymous):

f(x) = x2 + 7x + 8 (a) Write f(x) in the form k(x + t)2 + r. Fill in the obtained values of k, t, r.......... how would i graph this?

OpenStudy (anonymous):

you need to complete the square

OpenStudy (anonymous):

f(x)=x^2+7x+(7/2)^2-(7/2)^2+8 f(x)=x^2+7x+49/4-49/4+8 f(x)=(x+7/2)^2-17/4

OpenStudy (anonymous):

It should say write in the form of k(x-t)^2+r

OpenStudy (anonymous):

The vertex would be at (t,r), axis of symmetry at x=t